OpenStudy (anonymous):

need help, one question. giving medal. i need help with question E ONLY, all the other answers are already answered.

OpenStudy (anonymous):

Scenario 1: A. Graph y = –2x2 + 6. Describe what you see. There is the vertex at (0,6) and the upside down U-shaped parabola. B. –2x2 + 6 is what kind of polynomial? Classify the polynomial by degree and number of terms. it is 2nd degree and there is 2 terms, it is a Quadratic C. Next, graph y = (-1/10)x2 + 6. Describe what you see. How is what you observe here different from what you observed in part A? in part A- X went from -2 to 2 and in part B- x went from -7.5 to 7.5 D. If you wanted a really steep hill for your roller coaster, would you use the graph of y = –2x2 + 6 or y = (-1/10)x2 + 6? Why? What part of the expression affected the steepness? y = -x²/10 + 6 is the best graph, because you don't want to travel down too quickly when using the graph of y = -2x² + 6 E. Using the same reasoning from part D, create an equation that will be steeper than both equations, and explain your reasoning.

Miracrown (miracrown):

I would actually disagree with you on part D. The question wants a "really steep hill," so you *do* want to "travel down too quickly."

OpenStudy (anonymous):

i guess, im really not sure, i missed my lesson on this so im really confused.

Miracrown (miracrown):

So now, let's talk about part E. The difference between the two equations is that one of them has a "2" where the other one has a "1/10". What do you think the graph would look like if you put a different number there? Like maybe a 1?

OpenStudy (anonymous):

for D does this look better - D. If you wanted a really steep hill for your roller coaster, would you use the graph of y = –2x2 + 6 or y = (-1/10)x2 + 6? Why? What part of the expression affected the steepness? y = -2x² + 6 is the best graph, because y = -x²/10 + 6 would travel to slow. and for E yes i think its 1

OpenStudy (anonymous):

but for e it has to be steeper than both equations.

OpenStudy (anonymous):

@miracrown

OpenStudy (anonymous):

@dan815 can you help me?

OpenStudy (vishweshshrimali5):

@dominiquebee72 , you can check your answer for part (E) by drawing graphs for the equations.

OpenStudy (vishweshshrimali5):

Check out this : http://www.wolframalpha.com/input/?i=y+%3D+-x%5E2+%2B+6%2C+y+%3D+-2x%5E2+%2B6 The purple line represents : y = -2x^2 + 6 The blue line represents : y = -x^2 + 6

OpenStudy (vishweshshrimali5):

Which one is steeper ?